What is the equation of the line that contains the point (0, -1) and has a slope 3?
Found 2 solutions by Alan3354, Cromlix:
Answer by Alan3354(69443)

What is the equation of the line that contains the point (0, -1) and has a slope 3?
--------------
y = 3x - 1 is one equation.

Answer by Cromlix(4381)

Hi there,
Using the equation of a line:
y - b = m(x - a)
With m = 3 and (a,b) = (0, -1)
y - (-1) = 3(x - 0)
y + 1 = 3x
y = 3x - 1
Hope this helps :-)